OpenStudy (anonymous):

Solve for the variable. v - 8 = -7

OpenStudy (nikato):

hint: add 8 to both sides

OpenStudy (anonymous):

ok ill try

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

is it 15

OpenStudy (anonymous):

v - 8 = -7 v= -7 + 8 v= 1

OpenStudy (anonymous):

1? are u sure??

OpenStudy (nikato):

yea, hes right

OpenStudy (anonymous):

ok thanks y'all

OpenStudy (nikato):

u can always substitute -1 into v and check if the equation is true

OpenStudy (anonymous):

oh ok thanks!

OpenStudy (dan815):